Write an equation of the line containing the given point and perpendicular to the given live(5,9) 3x+y=4
Answer by mananth(16949)

3 x + 1 y = 4
Find the slope of this line
y = -3 x + 4
Compare this equation with y=mx+b
slope m = -3

The slope of a line perpendicular to the above line will be the negative reciprocal
m1*m2=-1

m= 1/3 ,point ( 5 , 9 )
Find b by plugging the values of m & the point in
y=mx+b
9 = 1/ 3*5 + b
b= 22/ 3
m= 1/3
The required equation is y = 1/ 3 x + 22/ 3
m.ananth@hotmail.ca